Update from Saturday 16th April. By Ian Robins and Paul Tucker.

Various work progressed on the 33's and Hymeks.

Ian writes

D6575 – Worked the DDEC as booked on Saturday.  The loco is due to have the brake blocks replaced
shortly, probably at BL on Saturday 30th April.

One of the automatic slack adjusters will also have to be changed and a serviceable spare is prepared and
ready.

D6566 - The Triple Pump has been fixed into position and wiring re-connected.
New silicon hose is being procured to replace the flexible sections in the cooling system pipework so once
the turbo and heat exchanger have been fitted we can plumb everything up, fill with water and see what
happens! The battery formerly installed in D6575 has arrived back following refurbishment by the supplier,
two defective cells were identified and replaced.
This set will now be installed in D6566 and once fitted we can then run the Triple Pump to get some water
pressure.  

Repairs to the main generator brush collector ring have been undertaken following advice from Bowers
Electricals. The damage is believed to have been caused in pre-preservation days by someone trying to
rotate the brush gear without having checked that the connecting braids were removed beforehand and
resulted in one of the collector ring tabs being bent upwards and the insulation cracking.   
So on Saturday the crack in the insulation was filled with a two part epoxy and once set the applicable
areas were given a liberal coat of anti-tracking varnish. On Sunday a second coat of ATV was applied and
the result can be seen in the attached pictures. Rest assured the piece of tissue put in to protect the
commutator that can be seen in the first picture was removed before the second picture was taken and
hasn’t been left inside the generator!  Bob C has made a bracket to lift the turbo.
Not to be caught out again like on D6575, the new turbo bellows has been fitted and the nuts wire locked
whilst the turbo is still out of the loco.
If you install the turbo prior to the bellows being fitted, the air filter box then has to be unbolted and moved
back.

D7017 – New oil pipes have been fitted to the transmission replacing the ones the ones that had split.  The
loco was started and it was confirmed that everything is now oil tight.

D7018 – The loco was started to asses the performance of the new cardan shaft between the engine and
transmission.  The verdict – now runs smooth as silk so £5k well spent!

D1010 – Julien performed further investigation into the reasons for lack of output from ‘A’ dynostarter.  The
likely cause seems to be dirty contacts so these have been cleaned and the situation will be re-assessed
next Saturday when a test run is planned.

Paul writes

This last weekend Leroy and I refitted D7017's engine water pump again as our engineer Bob was not
happy after changing it over Easter. He has spent at least 4 days working on it. This time it looks to be
better,as it is not leaking. Leroy and Gordon fitted the two new oil pipes from the Transmission heat
exchanger to the torque converter. It was discovered last weekend that these were leaking oil so new ones
were required. Bob meantime fitted the spacer to D7018's main drive shaft and during Saturday both
Hymeks were heard to run, although not at the same time. On Sunday I fitted 3 pipe clamps, to the water
system, and tried a 4th.


Pictures below from Ian showing turbo and repairs to main generator brush collecting ring.